OpenStudy (anonymous):

OMG I REALLY DONT UNDERSTAND... In parallelogram ABCD, angle A is 5x + 2 and angle B is 108° Plzzz explain to me

jimthompson5910 (jim_thompson5910):

in any parallelogram, the adjacent angles are supplementary this means that they add to 180 degrees

jimthompson5910 (jim_thompson5910):

So A+B = 180 (5x+2) + (108) = 180 5x + 2 + 108 = 180 .... .... x = ??

OpenStudy (anonymous):

14?

jimthompson5910 (jim_thompson5910):

and if x = 14, then A = 5x + 2 A = 5*14 + 2 A = ??

OpenStudy (anonymous):

72?

jimthompson5910 (jim_thompson5910):

good, so A = 72, B = 108 C = 72 (since C is opposite A) D = 108 (D is opposite B)

OpenStudy (anonymous):

thanks :)

jimthompson5910 (jim_thompson5910):

you're welcome